# LoRA-Pro: Are Low-Rank Adapters Properly Optimized?

### v1

@@ 5 Experimental Results
In this section, we evaluate our LoRA-Pro method across various natural language understanding datasets. To provide a comprehensive comparison, we include several baseline methods: 1) full fine-tuning and the standard LoRA [Hu et al., 2022]. 2) LoRA variants maintaining the original structure, such as rsLoRA [Kalajdzievski, 2023], LoRA+ [Hayou et al., 2024], PiSSA [Meng et al., 2024], 3) oRA variants with modified structures, including DoRA [Liu et al., 2024] and AdaLoRA [Zhang et al., 2023].
The results are shown in Table 1. We fine-tune the T5-base model [Raffel et al., 2020] with the baseline methods on a subset of GLUE datasets. From Table 1, we observe that LoRA-Pro achieves the highest scores on 3 out of 5 datasets and the highest average score across all 5 datasets. Moreover, on average over 5 datasets, LoRA-Pro suppass standard LoRA [Hu et al., 2022] with a margin of 6.72. These results validate the effectiveness of our methods.
Table 1: Results on fine-tuning T5-base with Full Fine-tuning and LoRA variants on a subset of GLUE datasets.
Method MNLI SST2 CoLA QNLI MRPC Avg.
@@
LoRA-Pro 86.92±0.08 94.46±0.24 82.25±1.01 92.89±0.12 87.50±0.65 88.80
@@ 6 Conclusion
In this paper, we introduce LoRA-Pro, a novel approach designed to bridge the performance gap between LoRA and full fine-tuning. To bridge the performance gap, we introduce the concept of Equivalent Gradient, which allows us to quantify the difference in the optimization process between LoRA and full fine-tuning. By minimizing this discrepancy, we derive the optimal closed-form updated solutions for LoRA. Moreover, we prove that the solutions guarantee the loss decease during optimization. These solutions not only apply a low-rank approximation to the fine-tuning matrix but also maintain consistency with the optimization of full fine-tuning, enabling more effective fine-tuning. Finally, we validate the effectiveness of our method through extensive experiments on natural language processing tasks.
Appendix A Notations
@@
$g=\frac{\partial L}{\partial W}\in\mathbb{R}^{m\times n}$ gradients of full rank fine-tuning
@@ Appendix B Proof of Theoretical Results
B.1 Proof of Theorem 3.3
Theorem B.1 Assume matrices $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are both full rank. For the objective $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, the solutions are given by: $g^{A}$ $=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA$ (15) $g^{B}$ $=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (16) Here, $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix.
Proof For simplicity, we denote $L=\|sBg^{A}+sg^{B}A-g\|_{F}^{2}$. To solve the optimization problem, we need to satisfy the following conditions: $\frac{\partial L}{\partial A}$ $=2sB^{T}(sBg^{A}+sg^{B}A-g)=0$ (17) $\frac{\partial L}{\partial B}$ $=2(sBg^{A}+sg^{B}A-g)sA^{T}=0$ (18) Given that matrices $A$ and $B$ are full-rank, $AA^{T}$ and $B^{T}B$ are invertible. And from Equation (18), we derive: $g^{B}=\frac{1}{s}gA^{T}(AA^{T})^{-1}-Bg^{A}A^{T}(AA^{T})^{-1}.$ (19) Substituting this into Equation (17), we obtain the following linear equation: $g^{A}[I-A^{T}(AA^{T})^{-1}A]=\frac{1}{s}(B^{T}B)^{-1}B^{T}g.$ (20) Here, we notice that the matrix $P=I-A^{T}(AA^{T})^{-1}A$ is a projection matrix with rank $n-r$. The solution to the linear equation (20) is: $g^{A}=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA,$ (21) where $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix. We take the solution (24) into Equation (19), we derive: $g^{B}=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX$ (22) While we have obtained closed-form solutions for $g^{A}$ and $g^{B}$, these solutions explicitly depend on the gradient of the matrix $W$, i.e., $g$, which is undesirable since $g$ is unknown during LoRA optimization. Fortunately, the solutions can be transformed into the forms of the gradients of standard LoRA, where the gradients are: $g^{A}_{lora}=sB^{T}g,\quad g^{B}_{lora}=sgA^{T}.$ (23) Therefore, the solutions to the optimization problem can be written as: $g^{A}$ $=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA,$ (24) $g^{B}$ $=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (25) In our method, we perform the standard forward and backward passes of LoRA, then adjust the gradients of A and B using Solutions (24) and (25), and subsequently update them.

### v2

@@ 3 Experimental Results
In this section, we present extensive experiments to evaluate the effectiveness of LoRA-Pro across various tasks and models. First, we assess natural language understanding capabilities using the GLUE benchmark by fine-tuning the T5-base (Raffel et al., 2020) model in Section 3.1. Next, we evaluate its capabilities in dialogue generation, mathematical reasoning, and code generation using the Llama-2-7B model (Touvron et al., 2023), covered in Section 3.2. We then examine LoRA-Pro’s effectiveness on image classification tasks using the CLIP-ViT-B/16 (Radford et al., 2021) model in Section 3.3. Finally, we conduct an ablation study of LoRA-Pro in Section 3.4.
Training details. To ensure a fair comparison, we align our experimental setup with that of LoRA-GA (Wang et al., 2024). By default, we fine-tune the model using the AdamW optimizer (Loshchilov & Hutter, 2019) with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay set to 0. We implement a cosine learning rate schedule with a warmup ratio of 0.03. LoRA is applied to all linear modules, excluding the embedding layer, normalization layer, and classification head. By default, we set the rank $r=8$ and $\alpha=16$.
For natural language understanding tasks, we fine-tune T5-base (Raffel et al., 2020) model with learning rate 1e-4. The sequence length is set to 128, and the training batch size is 32. For dialogue generation, mathematical reasoning and code generation tasks, we fine-tune the Llama-2-7B (Touvron et al., 2023) model with a learning rate of 2e-5. We set the sequence length to 1024 and the macro batch size to 32. For image classification tasks, we fine-tune the CLIP-ViT-B/16 (Radford et al., 2021) model with learning rate 1e-4. The classifier is obtained using prompts such as “a photo of a {class}” and kept frozen during fine-tuning. And the training batch size is set to 64.
All experiments are conducted on NVIDIA RTX A6000 GPUs. To obtain a reliable estimate of model performance, we perform three runs with different random seeds and report the average and standard deviation of the results.
3.1 Results on Natural Language Understanding Tasks
Table 1: Results of fine-tuning T5-base using full fine-tuning and various LoRA variants on a subset of the GLUE datasets. The LoRA rank is set to 8 by default. Bold and underline indicate the highest and second-highest scores, respectively.
Method MNLI SST2 CoLA QNLI MRPC Average
@@ 3.1 Results on Natural Language Understanding Tasks
LoRA-Pro 86.03±0.19 94.19±0.13 81.94±0.24 93.42±0.05 86.60±0.14 88.44
In this section, we evaluate our LoRA-Pro across various natural language understanding datasets. To provide a comprehensive comparison, we include several baseline methods: 1) full fine-tuning and the standard LoRA (Hu et al., 2022). 2) LoRA variants maintaining the original structure, such as rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and LoRA-GA (Wang et al., 2024), 3) LoRA variants with modified structures, including DoRA (Liu et al., 2024) and AdaLoRA (Zhang et al., 2023).
The results are presented in Table 1. We fine-tune the T5-base model (Raffel et al., 2020) with the baseline methods on a subset of GLUE datasets: MNLI, SST2, CoLA, QNLI, and MRPC. As shown in Table 1, we observe that LoRA-Pro achieves the highest scores on 3 out of 5 datasets and the highest average score across all 5 datasets, and achieves the highest accuracy on average over the 5 datasets. Specifically, on average over 5 datasets, LoRA-Pro surpasses standard LoRA (Hu et al., 2022) with a margin of 6.36. And LoRA-Pro even achieve higher than full fine-tuning. This superior performance may be attributed to overfitting in full fine-tuning, where optimizing all model parameters can lead to overfitting on the training data, thus reducing the model’s generalization to the test set. This effect is particularly pronounced on small datasets, such as MRPC, which contains only 3.7k training data. These results validate the effectiveness of our methods.
Table 2: Fine-tuning results of Llama-2-7B model. We fine-tune the Llama-2-7B model using full fine-tuning and LoRA variants on subsets of the WizardLM (Xu et al., 2024), MetaMathQA (Yu et al., 2024), and CodeFeedback (Zheng et al., 2024b) datasets, respectively. And we assess dialogue generation, mathematical reasoning, and coding abilities on MT-Bench, GSM8K, and HumanEval datasets. Bold and underline indicate the highest and second-highest scores, respectively.
MT-Bench GSM8K HumanEval
Full FT 5.30±0.11 59.36±0.85 35.31±2.13
LoRA 5.61±0.10 42.08±0.04 14.76±0.17
PiSSA 5.30±0.02 44.54±0.27 16.02±0.78
rsLoRA 5.25±0.03 45.62±0.10 16.01±0.79
LoRA+ 5.71±0.08 52.11±0.62 18.17±0.52
DoRA 5.97±0.02 53.07±0.75 19.75±0.41
AdaLoRA 5.57±0.05 50.72±1.39 17.80±0.44
LoRA-GA 5.95±0.16 53.60±0.30 19.81±1.46
LoRA-GA (rank=32) 5.79±0.09 55.12±0.30 20.18±0.19
LoRA-GA (rank=128) 6.13±0.07 55.07±0.18 23.05±0.37
LoRA-Pro 5.86±0.06 54.23±0.79 22.76±0.35
LoRA-Pro (rank=32) 6.01±0.05 55.14±1.73 28.05±0.00
LoRA-Pro (rank=128) 5.68±0.14 56.48±0.23 34.55±2.46
3.2 Results on Large Language Models
In this section, we evaluate the performance of LoRA-Pro on large language models, focusing on dialogue generation, mathematical reasoning, and code generation capabilities. Our experimental setup follows the configuration used in LoRA-GA (Wang et al., 2024).
  • For the dialogue generation task, we fine-tune the Llama-2-7B (Touvron et al., 2023) model on a 52k subset of the WizardLM dataset (Xu et al., 2024) and evaluate it using the MT-Bench dataset (Zheng et al., 2024a). GPT-4 is used to assess the quality of the model’s responses, and we report the first-turn score as the metric.
  • For the math task, we fine-tune the Llama-2-7B (Touvron et al., 2023) model on a 100k sample from the MetaMathQA dataset (Yu et al., 2024). The model is then evaluated on the GSM8K test set (Cobbe et al., 2021), and we report the accuracy as the metric.
  • For the coding task, we fine-tune the Llama-2-7B (Touvron et al., 2023) model on a 100k subset of the CodeFeedback dataset (Zheng et al., 2024b) and test it on the HumanEval dataset (Chen et al., 2021), reporting the PASS@1 metric.
We compare LoRA-Pro with several baselines, including full fine-tuning, LoRA (Hu et al., 2022), PiSSA (Meng et al., 2024), rsLoRA (Kalajdzievski, 2023), LoRA+(Hayou et al., 2024), DoRA(Liu et al., 2024), AdaLoRA (Zhang et al., 2023), and LoRA-GA (Wang et al., 2024). By default, we set the rank to 8 and $\alpha=16$. Following LoRA-GA (Wang et al., 2024), we initialize the scaling factor as in rsLoRA (Kalajdzievski, 2023), i.e., $s=\frac{\alpha}{\sqrt{r}}$.
Table 2 presents our experimental results, which demonstrate LoRA-Pro’s superior performance. With a rank of 8, LoRA-Pro achieves notable improvements over the original LoRA: 0.25 on MT-Bench, 11.98 on GSM8K, and 8.00 on HumanEval. When compared to the second-best PEFT method, LoRA-GA, LoRA-Pro shows consistent gains: 0.46 on GSM8K and a substantial 2.95 on HumanEval. These results validate the effectiveness of our LoRA-Pro method.
Interestingly, we observe that full fine-tuning unexpectedly underperforms on MT-Bench. We attribute this to potential discrepancies between the WizardLM training data distribution and the MT-Bench evaluation set. The extensive learning capacity of full fine-tuning may lead to overfitting on the training distribution, compromising generalization to MT-Bench. Since LoRA-Pro aligns more closely with full fine-tuning during optimization, its relatively poor performance on MT-Bench may also be attributed to overfitting.
To further explore the scalability of our method, we increase the rank in LoRA-Pro from 8 to 128. Our observations reveal a clear trend: as the rank increases, the performance gap between LoRA-Pro and full fine-tuning narrows rapidly. Notably, LoRA-Pro consistently outperforms LoRA-GA at the same ranks on both GSM8K and HumanEval datasets. At rank 32, LoRA-Pro surpasses LoRA-GA by 0.02 on GSM8K and 7.87 on HumanEval. This performance disparity becomes even more pronounced at rank 128, where LoRA-Pro outperforms LoRA-GA by 1.41 on GSM8K and 11.5 on HumanEval. These results demonstrate the superior scalability and effectiveness of LoRA-Pro across various rank settings.
3.3 Results on Image Classification Tasks
Table 3: Fine-tuning results of CLIP-ViT-B/16 on image classification tasks. We fine-tune CLIP-ViT-B/16 using full fine-tuning and LoRA variants across StanfordCars, DTD, EuroSAT, GTSRB, RESISC45, SUN397, and SVHN datasets. Bold indicates the highest results, while underline represents the second-highest results.
Method Cars DTD EuroSAT GTSRB RESISC45 SUN397 SVHN Average
Zero-shot 63.75 44.39 42.22 35.22 56.46 62.56 15.53 45.73
Full FT 84.23±0.06 77.44±0.19 98.09±0.03 94.31±0.28 93.95±0.0 75.35±0.10 93.04±0.18 88.06
LoRA 72.81±0.13 73.92±0.38 96.93±0.07 92.40±0.10 90.03±0.14 70.12±0.18 88.02±0.07 83.46
rsLoRA 82.38±0.20 78.03±0.76 98.06±0.08 95.04±0.11 93.96±0.18 75.38±0.24 92.74±0.18 87.94
LoRA+ 72.87±0.18 74.07±0.45 97.01±0.02 92.42±0.18 89.96±0.11 70.17±0.15 88.08±0.05 83.51
DoRA 73.72±0.06 73.72±0.33 96.95±0.01 92.38±0.17 90.03±0.08 70.20±0.19 88.23±0.05 83.48
LoRA-Pro 85.87±0.08 78.64±0.25 98.46±0.03 95.66±0.05 94.75±0.21 76.42±0.14 94.63±0.20 89.20
In this section, we assess the performance of LoRA-Pro on image classification tasks. To provide a comprehensive comparison, we compare it with several baselines: zero-shot CLIP (Radford et al., 2021), full fine-tuning, vanilla LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), and DoRA (Liu et al., 2024).
We fine-tune the CLIP-ViT-B/16 (Radford et al., 2021) model on various datasets, including StanfordCars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Houben et al., 2013), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2010), and SVHN (Netzer et al., 2011). Accuracy is used as the evaluation metric. During fine-tuning, only the visual backbone of the CLIP-ViT-B/16 model is updated, while the classifier, derived from prompts such as “a photo of a {class}”, remains frozen.
The results are presented on Table 3. LoRA-Pro achieves the highest accuracy across all seven datasets. Specifically, on average, LoRA-Pro surpasses zero-shot classification by 43.47, outperforms LoRA (Hu et al., 2022) by 5.74, and exceeds rsLoRA (Kalajdzievski, 2023) by 1.26. These results validate the effectiveness of our LoRA-Pro method.
3.4 Ablation Study
Table 4: Ablation study on the selection of different X in LoRA-Pro.
choice of X MT-Bench GSM8K HumanEval
Zero 5.76±0.02 53.83±1.16 22.96±1.96
Sylvester (Thm. 2.3) 5.86±0.06 54.23±0.79 22.76±0.35
Symmetry (Eq. (16)) 5.63±0.12 54.46±0.88 22.56±1.06
Ablation study on the selection of X. Based on Theorem 2.1, in LoRA-Pro, the matrix $X$ can be chosen arbitrarily. While its selection does not affect the equivalent gradient, it does influence the updates of matrices $A$ and $B$ in LoRA. Here, we conduct an ablation study on the choice of $X$.
We compare three possible values for $X$. 1) Zero solution: In this simplest case, we set $X=\mathbf{0}$. 2) Sylvester solution: Here, $X$ is obtained by solving the Sylvester equation, as described in Theorem 2.3. 3) Symmetry solution: This approach aims to balance the contributions of both terms in the equation $\tilde{g}=sg^{B}A+sBg^{A}$, enforcing the condition $g^{B}A=Bg^{A}$. For the symmetry solution, solving for $X$ yields:
$X=-\frac{1}{2s}B(B^{T}B)^{-1}B^{T}gA(A^{T}A)^{-1}A=-\frac{1}{2s^{2}}B(B^{T}B)^{-1}B^{T}g^{B}_{lora}(A^{T}A)^{-1}A.$ (16)
The comparison of the selection of $X$ is presented in Table 4. As shown in the table, the results obtained from different choices of $X$ are similar, and, except on MT-Bench, they outperform the other PEFT baselines listed in Table 2. However, considering the high variance in the results for both the zero solution and the symmetry solution, we ultimately select the sylvester solution as the default choice for the matrix $X$.
Ablation study on the full-rank assumption. In Theorem 2.1, we assume that the matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ are full-rank during training. Our goal here is to verify whether this assumption holds in practice. We track the rank changes of all $A$ and $B$ matrices during the fine-tuning process of Llama-2-7B on the MetaMathQA (Yu et al., 2024) dataset.
In Figure 2, we illustrate the rank changes of matrices $A$ and $B$ from the $q$ projection of layer 9 during the training process, with rank set to 8 and 32, resipectively. We observed that, although $A$ and $B$ do not initially satisfy the full rank assumption (with matrix $B$ initialized as a zero matrix), both matrices achieve full rank after the first update step. The rank behavior of $A$ and $B$ in other layers also exhibits similar results.
Figure 2: Visualization of matrix ranks of A and B during training, with ranks set to $8$ and $32$, respectively.
This observation provides practical evidence that the assumption in Theorem 2.1 is reasonable and supports the validity of the proposed solutions.
Table 5: We compare LoRA, LoRA-Pro, and Full Fine-Tuning in terms of memory cost, training time, and performance on the MT-Bench, GSM8K, and HumanEval datasets. Memory cost is measured using a single A6000 GPU with a batch size of 1. Training time is recorded on the WizardLM dataset using 8 A100 GPUs with DeepSpeed ZeRO-2 stage optimization.
Memory Cost Training Time MT-Bench GSM8K HumanEval
Full FT $>$ 48 GB 2h 33min 5.30±0.11 59.36±0.85 35.31±2.13
LoRA 22.26 GB 1h 22min 5.61±0.10 42.08±0.04 14.76±0.17
LoRA-Pro 23.05 GB 1h 23min 5.86±0.06 54.23±0.79 22.76±0.35
Memory footprint and training time. Here, we evaluate the additional costs associated with using LoRA-Pro compared to LoRA. We primarily focus on comparing the differences in memory cost and training time between LoRA-Pro, LoRA, and full fine-tuning. The results of the experiments are presented in Table 5. We measure memory cost using a single A6000 GPU with a batch size of 1. Training time is recorded on the WizardLM dataset using 8 A100 GPUs with DeepSpeed (Rasley et al., 2020) ZeRO-2 stage optimization.
The results are shown in Table 5. From the table, we observe the following: 1) LoRA-Pro requires approximately 0.8GB more GPU memory compared to LoRA. This difference likely stems from the need to compute $B^{T}B$, $AA^{T}$, and their inverses during the calculation of the optimal gradients. 2) Surprisingly, the training time for LoRA-Pro is nearly identical to that of LoRA, with only about a 1% increase in additional training time. We attribute this to the fact that the matrices $A$ and $B$ in LoRA are low-rank. Consequently, the extra computations required by LoRA-Pro (such as matrix inversion and the calculation of $X$) are performed on small $r\times r$ matrices, resulting in negligible computational overhead.
Considering that LoRA-Pro uses less memory and trains faster than full fine-tuning, while also providing performance improvements over LoRA, we believe that the additional memory and training time costs are acceptable.
4 Related Work
Parameter-Efficient Fine-Tuning. Given the huge size of foundation models, recent research has focused on developing parameter-efficient fine-tuning methods (Hu et al., 2022; Liu et al., 2024; Ding et al., 2023; Houlsby et al., 2019; Liu et al., 2023; Lester et al., 2021). These methods aim to reduce the cost of fine-tuning by adjusting only a small portion of the model’s parameters. Generally, these methods fall into two main categories. The first category is adapter tuning (Houlsby et al., 2019; Sung et al., 2022; He et al., 2021; Zhang et al., 2024; Bapna & Firat, 2019; Hu et al., 2022), which involves inserting small neural network modules, called adapters, into specific layers of the model. During fine-tuning, we keep the model frozen and only fine-tune the lightweight adapter modules, significantly reducing the memory footprint for fine-tuning. The second category is prompt tuning (Lester et al., 2021; Zhou et al., 2022; Li & Liang, 2021; Liu et al., 2022). Prompt tuning adapts the models to specific tasks by adding specially designed prompts or learnable tokens to the input data, rather than directly modifying the internal parameters of foundation models. In this paper, we focus on LoRA (Hu et al., 2022), a prominent method within the realm of adapter tuning.
Low-Rank Adaptation. Low-rank adaptation, initially referred to as LoRA (Hu et al., 2022), has evolved into a broad category encompassing parameter-efficient fine-tuning methods based on low-rank approximations (Hu et al., 2022; Liu et al., 2024; Hayou et al., 2024; Kalajdzievski, 2023; Zhang et al., 2023; Kopiczko et al., 2024; Hyeon-Woo et al., 2022; Zhang & Pilanci, 2024; Wang et al., 2024; Zhao et al., 2024; Wang et al., 2024). LoRA (Hu et al., 2022) assumes that the changes in the weights of pre-trained models exhibit a low-rank structure. Consequently, it re-parameterizes these changes as the product of low-rank matrices, thereby reducing the cost during fine-tuning.
Several variants of LoRA have been proposed to address different aspects of this approach. For example, DoRA (Liu et al., 2024) improves LoRA (Hu et al., 2022) by incorporating a learnable magnitude vector to re-scale the normalized product of low-rank matrices. Another variant, rsLoRA (Kalajdzievski, 2023), introduces a new scaling factor to stabilize training in high-rank scenarios. LoRA+(Hayou et al., 2024) improves upon LoRA by applying different learning rates to the two low-rank matrices. Additionally, Galore (Zhao et al., 2024) employs SVD to project the gradients and its first and second momentum of full training into a low-rank space, thereby reducing the memory footprint during pre-training and fine-tuning.
@@ 5 Conclusion
In this paper, we introduce LoRA-Pro, a novel approach designed to bridge the performance gap between LoRA and full fine-tuning. We have discovered that using LoRA for fine-tuning is equivalent to fine-tuning the original weights with a virtual equivalent low-rank gradient. Based on this insight, we propose adjusting the gradients of matrices $A$ and $B$ to make the equivalent gradient match the true full fine-tuning gradient, thereby reducing their performance gap. Fortunately, we theoretically prove that there exists an optimal closed-form solution for updating matrices $A$ and $B$, which are applied during fine-tuning in LoRA-Pro. To validate the effectiveness of our method, we conduct extensive experiments across various domains, including natural language understanding, dialogue generation, mathematical reasoning, code generation, and image classification tasks. The results demonstrate that LoRA-Pro significantly improves LoRA performance and narrows the performance gap with full fine-tuning.
Appendix A Notation
@@
$g=\frac{\partial L}{\partial W}\in\mathbb{R}^{m\times n}$ gradients of full fine-tuning
@@ Appendix B Proof of Theoretical Results
B.1 Proof that the Equivalent Gradient is Low-Rank
Lemma.
Assume $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ and $g^{B}\in\mathbb{R}^{m\times r},g^{A}\in\mathbb{R}^{r\times n}$ represent matrices and their corresponding gradients in LoRA optimization. We demonstrate that the equivalent gradient: $\tilde{g}=sg^{B}A+sBg^{A},$ (17) where $s>0$ is the scaling factor, has matrix rank at most $2r$.
Proof.
Since matrix rank satisfies the property of subadditivity, we have:
$rank(\tilde{g})=rank(sg^{B}A+sBg^{A})\leq rank(g^{B}A)+rank(Bg^{A}).$ (18)
Furthermore, for any matrices $A$ and $B$, $rank(AB)\leq\min(rank(A),rank(B))$. Therefore, we can bound the ranks as follows:
$rank(g^{B}A)$ $\leq\min(rank(g^{B}),rank(A))\leq r$ (19)
$rank(Bg^{A})$ $\leq\min(rank(B),rank(g^{A}))\leq r$ (20)
Thus, in conclusion, the equivalent gradient has a rank of at most $2r$:
$rank(\tilde{g})$ $\leq rank(g^{B}A)+rank(Bg^{A})$ (21)
$\leq\min(rank(g^{B}),rank(A))+\min(rank(B),rank(g^{A}))$ (22)
$\leq 2r.$ (23)